OpenStudy (anonymous):

Cesar needs 1 1/2 cups of flour to make a cake that will serve 14 people. If Cesar wants to make enough cake to serve 28 people, how many cups of flour will he need? 2 1/2 2 6 3 1/4

OpenStudy (anonymous):

im stuck between 2 and 2 1/2 because it does say 1 1/2 will serve 14 people and 14 + 14=28 meaning adding 1 1/2 + 1 1/2 =2.1

OpenStudy (anonymous):

1 1/2=1.5 so answer should be 1.5+1.5=3 but it's not in the answer i think there is some problem with the answer